OpenStudy (anonymous):

Among all pairs of numbers whose difference is 10, find a pair whose product is as small as possible. What is the minimum product? a. the pairs of numbers whose difference is 10 and whose product is as small as possible is___ b. The minimum product is___

OpenStudy (ranga):

Let one number be x. The other number can be (x + 10) to give a difference of 10. Product P = x(x + 10) = x^2 + 10x For minimum P, find dP/dx = 0 and solve for x.

OpenStudy (texaschic101):

11 - 1 = 10 11 * 1 = 11 I can't seem to find anything smaller

OpenStudy (ranga):

The numbers can be positive or negative. P = x^2 + 10x dP/dx = 2x + 10 = 0 x = -5 One number is -5 and the other number is +5 Their product is -25 and that is the minimum.

OpenStudy (texaschic101):

oh.....5 - (-5) = 10 5 * -5 = -25 I didn't think of that

OpenStudy (ranga):

You can try any other combination and the product will always be greater than -25. If they place a restriction that the numbers cannot be negative then the obvious choice is 0 & 10.

OpenStudy (texaschic101):

I get it now.....thumbs up to you ranga :)

OpenStudy (ranga):

thank you.